# Automated Model Tuning for Multifidelity Uncertainty Propagation in Trajectory Simulation

### v1

@@ 3 Application
This work demonstrates the practical aspects of automated model tuning for multifidelity UQ in the context of trajectory simulation for entry, descent, and landing (EDL). Onboard trajectory simulation is emblematic of problems where little a priori knowledge about optimal low-fidelity model settings and model statistics (e.g., correlations) may be available and computational efficiency is critical. In order to be practically useful then, multifidelity UQ must be performed in an online setting where upfront model tuning and pilot sampling costs cannot be ignored as they often are in existing literature.
The particular problem considered here is the trajectory simulation of a flight test conducted on an atmospheric entry system called the Adaptable Deployable Entry and Placement Technology (ADEPT) [26]. ADEPT is an umbrella-like heatshield made of 3-D woven carbon fabric that acts as both a thermal protection system and a structural membrane. In smaller sizes, ADEPT can be equipped on small satellites for direct entry and/or aerocapture. At larger sizes it could enable human-scale missions with a large drag area with reduced launch sizes and weights. The ADEPT Sounding Rocket One (ADEPT SR-1) flight test was performed to test its ability to deploy exo-atmospherically into the specified aerodynamic configuration as well as to evaluate the aerodynamics of its first-of-a-kind faceted geometry (See Figure 2). Likewise, the novelty of this test flight has led to ADEPT SR-1 becoming a testbed for trajectory simulation and UQ techniques.
@@
Ultimately, the trajectory simulation gives a means for predicting 15 quantities of interest of the flight given a sample of its more than 80 uncertain input parameters (e.g., deployment orientation, atmospheric conditions, etc.). Though the MC method was shown to produce reasonable results, the analysis was revisited with multifidelity UQ to show increased computational efficiency [3]. Four models were used in that work: (1) the high-fidelity POST2 model used in the original MC-based work, (2) a POST2 model with reduced physics, i.e., simplified atmospheric model, (3) a POST2 model with coarse timestep, i.e., reduced integration accuracy, and (4) a machine learning surrogate model where a support vector machine was trained on samples of the high-fidelity model. The models had varying levels of computational speedup compared to the high-fidelity model, as seen in Table 2. Here, it is assumed that the training of the machine learning surrogate model (including generation of training data) is an offline cost and is thus not factored into the speedup value. Notably, the reduced physics and coarse timestep also had hyperparameters ($\Delta t$) which were hand-selected with best engineering judgment. This work focuses on exploring the benefits of automated model tuning of the time step hyperparameters relative to the performance of using hand tuned values.
Table 2: Models considered for multifidelity UQ.
@@
The proposed approach was implemented using Sandia National Laboratories’ UQ code, Dakota. Figure 3 shows a diagram illustrating the implementation approach. Dakota controls the solution of the bi-level model tuning optimization problem (Equation (9)), using global surrogate-based optimization (i.e., the EGO algorithm) on the outer loop, passing values of $\underline{\beta}$ to ACV sample allocation optimization at each iteration. To solve the ACV problem, calls are dispatched to high- and low-fidelity models for pilot sampling to estimate the required model statistics for calculating estimator variance.
@@ 4 Results
This section presents results from two numerical examples that apply model tuning for multifidelity UQ in the context of the ADEPT application. The first example focuses on tuning the time step of the Coarse Time Step model (Table 2) only. In this simpler 1D setting, the end-to-end online multifidelity UQ analysis depicted in Figure 1 is carried out to evaluate the performance of the approach in a realistic setting where pilot sampling costs and errors are accounted for in the construction of multifidelity estimators.
@@
    Hand Selected - multifidelity UQ estimators generated with fixed models borrowed from previous work [3] that were selected with engineering judgement (Table 2). For example 1, online pilot solution mode will be used to estimate model statistics on the fly during the sample allocation optimization.
@@
    Best Case - multifidelity UQ estimators generated with a priori knowledge of optimal hyperparameter values (no tuning cost) and exact model statistics (offline pilot solution mode with $N_{pilot}=10000$).
@@
Figure 5: Variance reduction as a function of low-fidelity model time steps for a) time of flight for example 1 (with fixed reduced physics model time step), b) time of flight for example 2, c) landing latitude for example 2 , and d) angle of attack @ 10km for example 2.
Figure 4 illustrates the effect of the tunable time step parameters on the low-fidelity models’ computational cost and correlations versus the other models for the time of flight QoI. While the cost of both the Reduced Physics and Coarse Time Step models decrease consistently with increasing time step as expected, the model correlations exhibit interesting and non-intuitive behavior. The correlations of both tunable models with respect to the High Fidelity and Machine Learning models exhibit relatively constant correlation over a large range of increasing time step, but show a drastic decrease/increase with respect to the High Fidelity/Machine Learning around $\Delta t=10^{-1}$. The complex nature of this cost versus correlation trade-off makes it difficult to make a priori assumptions about valid hand tuned time step values. Note that ACV methods also take advantage of the correlations amongst the low-fidelity models themselves, further complicating manual hyperparameter tuning.
@@
that is possible with the ACV-MF estimator as a function of the tunable time step parameters for a computational budget of 1000 equivalent high-fidelity model evaluations. Here, the estimator variance was evaluated by discretizing each time step on a grid of 25 equally (log) spaced points using offline pilot solution mode with $N_{pilot}=10k$ for a reference solution. Example 1 is a 1D model tuning problem focused on estimating the time of flight QoI (Figure 5 (a)) while example 2 is a 2D model tuning problem considering the time of flight, landing latitude, and angle of attack at 10km QoIs (Figure 5 (b)-(d)). These plots help to illustrate the nature of the model tuning optimization problem that must be solved, effectively providing a visualization of the objective function to be minimized with EGO. The plots also show the time step values for the Hand Selected and Best Case solutions, highlighting the large difference between those settings as well as the resulting variance reduction. In particular, the optimal time step setting for the Reduced Physics model is significantly larger than the hand selected value for all three QoIs in Figures 5 b-d. This indicates that there may be a tendency by practitioners to overestimate how accurate a low-fidelity model must be to be useful within a multifidelity framework, where models with large absolute errors can still be valuable provided they are adequately correlated.
@@ 4.1 Example 1 - Single Tunable Hyperparameter
Example 1 carries out the full online ACV estimator with the model tuning procedure shown in Figure 1, using the bilevel optimization strategy in Equation (9) to select an optimal timestep for the Coarse Time Step model. Estimators for the time of flight QoI are formed using different numbers of pilot samples, $N_{pilot}=\{10,50,100\}$, and EGO iterations, $N_{iter}=\{5,10,20\}$ [^5] EGO uses an initial sample size of 3 to generate a surrogate model, so these values represent 2, 7, and 17 EGO iterations beyond the initial sample, respectively, to solve the model tuning optimization problem. One hundred random trials are performed for each model tuning estimator as well as the Monte Carlo, Hand Selected, and Best Case solution types described above for three different computational budgets, $C_{budget}=\{500,1000,2000\}$, defined in terms of equivalent high-fidelity model cost. Note that the Best Case solution uses an optimal time step value of $\Delta t=0.065\;\mathrm{sec}$ and the Hand Selected case uses $\Delta t=0.1\;\mathrm{sec}$ (see Figure 5 a). Each solution approach is compared in terms of the MSE defined in Equation (12) with a reference solution computed using Monte Carlo with $N=100k$ samples. For all ACV solutions, the ACV-MF estimator (Table 1) is used.
Figure 6: Model tuning optimization results across 100 random trials for example 1. Distributions of the resulting optimal time step values are shown versus the estimator variance curve being minimized for a) different numbers of pilot samples with fixed EGO iterations = 5, and b) different numbers of EGO iterations with fixed number of pilot samples = 25.
@@
Results for the solution of the model tuning optimization problem (Equation (9)) are shown in Figures 6 and 7. Figure 6 shows the optimal time step values returned by the EGO algorithm across the 100 random trials for different numbers of $N_{pilot}$ and $N_{iter}$, represented as distributions computed with kernel density estimation. Compared with the relative estimator variance curve that is being optimized, it can be seen in (a) that the EGO-optimized time steps noticeably improve with increasing $N_{pilot}$. In particular, for $N_{pilot}=10$, it can be seen that a relatively large percentage of optimized time steps are larger values that are sub-optimal in the sense that the estimator variance is high. For increasing $N_{iter}$ in (b), the trend is less clear, indicating that $N_{iter}=5$ seems to be sufficient and there are diminishing returns beyond this value. More detailed optimization results can be found in Appendix B and Figure 17. It is worth noting that while relative estimator variance objective appears simple to optimize in Figure 6, this curve was generated using $N_{pilot}=50k$ and is particularly noisy for the smaller values of $N_{pilot}$ used for model tuning, making the model tuning optimization problem challenging even in one dimension. This challenge is expanded upon in Appendix C and Figure 18.
@@
Figure 7: The overhead from solving the model tuning optimization problem as a percentage of overall computational budget for example 1. Box plots indicating the distribution of overheads for 100 random trials are shown for a) different numbers of pilot samples with fixed EGO iterations = 5, and b) different numbers of EGO iterations with fixed number of pilot samples = 25.
@@
Figure 7 presents the computational cost of the model tuning optimization problem for different numbers of $N_{pilot}$ and $N_{iter}$. Here, the model tuning overhead is represented as a percentage of the total budget, $100\times\frac{C_{\mathcal{T}}}{C_{budget}}$, for different total budgets $C_{budget}=\{500,1000,2000\}$. Since there is variability in the progression of the EGO algorithm, the overheads are shown as box plots illustrating the variation across the 100 random trials. Intuitively, this shows that the overhead incurred by model tuning prior to multifidelity UQ is less significant with increasing total computational budget. So for $C_{budget}=500$, using $N_{pilot}=100$ or $N_{iter}=20$ depletes a large percentage of the budget (between $10-30\%$), while these percentages drop below $5\%$ for $C_{budget}=2000$. Table 3 summarizes the performance obtained through model tuning optimization along with the incurred overhead cost. Here, the $\%$ of random model tuning trials that returned hyperparameters that result in lower estimator variance than the Hand Selected solution are reported along with the average tuning cost (in terms of number of high-fidelity simulations). It can be seen that performance increases with increasing $N_{pilot}$, albeit with diminishing returns, while performance does not increase with increasing $N_{iter}$, despite dedicating increased cost for model tuning. This result indicates that it is more worthwhile to consider larger values for $N_{pilot}$ rather than $N_{iter}$ for this one-dimensional tuning problem.
@@
Figure 8: MSE versus budget for model tuning with 50 pilot samples and 5 EGO iterations compared with Monte Carlo, ACVs with hand-selected models, and the best-case ACV solution.
@@
Figure 9: MSE versus number of pilot samples compared with the Monte Carlo, hand-selected, and best case estimators for computational budgets of 500 (left), 1000 (middle), and 2000 (right) high-fidelity simulations..
Figure 10: MSE versus number of EGO iterations compared with the Monte Carlo, hand-selected, and best case estimators for computational budgets of 500 (left), 1000 (middle), and 2000 (right) high-fidelity simulations..
@@ 4.2 Example 2 - Tuning of Multiple Hyperparameters
Example 2 studies the tuning of time steps for both the Reduced Physics, $\Delta t_{RP}$, and Coarse Time Step, $\Delta t_{CST}$, models through the solution of Equation (9). Similar to Example 1, model tuning performance is evaluated for different numbers of pilot samples, $N_{pilot}=\{10,50,100\}$, and EGO iterations, $N_{iter}=\{10,20,30\}$ [^6] EGO uses an initial sample size of six to generate a surrogate model in two dimensions, so these values represent 4, 14, and 24 EGO iterations beyond the initial sample, respectively. Performance was assessed in terms of estimator variance, rather than the more computationally-intensive MSE considered in Example 1, allowing for the consideration of additional QoIs (the landing latitude and angle of attack at 10km) and larger computational budgets. The variance reduction contours for each of the three QoIs can be seen in Figure 5 (b)-(d), showing the varied nature of the model tuning optimization problems (perhaps simpler for landing latitude relative to time of flight and angle of attack). A maximum computational budget of $C_{budget}=10000$ is considered to better understand the relative advantages of model tuning for more realistic analyses [^7] The original Monte Carlo analysis for ADEPT used $C_{budget}=8000$ high-fidelity simulations [27]. Note that while the model tuning procedure is simply repeated for each QoI individually for the results presented next, the approach could be extended to find hyperparameters that perform optimally for the joint estimation of all QoIs simultaneously using multioutput ACVs [33].
Figure 11: Model tuning optimization results for different numbers of pilot samples and fixed EGO iterations = 10 across 100 random trials for a) time of flight, b) landing latitude, and c) angle of attack QoIs.
Figure 12: The overhead from solving the model tuning optimization problem as a percentage of overall computational budget for example 2. Box plots indicating the distribution of overheads for 100 random trials are shown for a) different numbers of pilot samples with fixed EGO iterations = 10, and b) different numbers of EGO iterations with fixed number of pilot samples = 50.
@@
The performance of the model tuning solutions was assessed with respect to the other baseline approaches in terms of variance reduction with respect to the Monte Carlo solution (Equation 13). In order to reduce the computation time needed to estimate the variance reduction, an interpolator was constructed to approximate $\mathbb{V}ar\!\left[\tilde{Q}\right$]
@@
Here, $\tilde{\mathbb{V}}$ is constructed from estimator variance data on a $25\times 25$ grid of equally (log) spaced time steps using offline pilot solution mode with $N_{pilot}=10k$ for a reference solution (i.e., the same data used to produce the contour plots in Figure 5). For the Hand Selected solution, $\tilde{\mathbb{V}}$ was evaluated with $\Delta t_{RP}=0.001$ and $\Delta t_{CST}=0.1$ while the Best Case solutions used different optimal time step settings for the different QoIs considered (see Figure 5). For the model tuning solution, the estimated variance from the interpolator was scaled to account for the computational overhead of solving Equation (9),
@@
Figure 13: Model tuning performance for different computational budgets using different numbers of pilot samples with fixed EGO iterations = 10. The box plots indicate the variance reduction achieved across 100 random trials for a) time of flight, b) landing latitude, and c) angle of attack QoIs, compared to variance reduction using hand-selected models and the best case variance reduction possible.
Figure 14: Model tuning performance for different computational budgets using different numbers of EGO iterations with fixed number of pilot samples = 50. The box plots indicate the variance reduction achieved across 100 random trials for a) time of flight, b) landing latitude, and c) angle of attack QoIs, compared to variance reduction using hand-selected models and the best case variance reduction possible.
@@
Figure 15: Model tuning performance when combined with parametric ACV search using the generalized multifidelity (GMF) sampling strategy. The box plots indicate the variance reduction achieved across 100 random trials for a) time of flight, b) landing latitude, and c) angle of attack QoIs, for a fixed computational budget of 5000 and EGO iterations = 10.

### v2

@@ 3 Application
This work demonstrates the practical aspects of automated model tuning for multifidelity uncertainty propagation in the context of trajectory simulation for entry, descent, and landing (EDL). Onboard trajectory simulation is emblematic of problems where little a priori knowledge about optimal low-fidelity model settings and model statistics (e.g., correlations) may be available and computational efficiency is critical. In order to be practically useful then, multifidelity estimators must be constructed in an online setting where upfront model tuning and pilot sampling costs cannot be ignored as they often are in existing literature.
The particular problem considered here is the trajectory simulation of a flight test conducted on an atmospheric entry system called the Adaptable Deployable Entry and Placement Technology (ADEPT) [30]. ADEPT is an umbrella-like heatshield made of 3-D woven carbon fabric that acts as both a thermal protection system and a structural membrane. In smaller sizes, ADEPT can be equipped on small satellites for direct entry and/or aerocapture. At larger sizes it could enable human-scale missions with a large drag area with reduced launch sizes and weights. The ADEPT Sounding Rocket One (ADEPT SR-1) flight test was performed to test its ability to deploy exo-atmospherically into the specified aerodynamic configuration as well as to evaluate the aerodynamics of its first-of-a-kind faceted geometry (See Figure 2). Likewise, the novelty of this test flight has led to ADEPT SR-1 becoming a testbed for trajectory simulation and uncertainty propagation techniques.
@@
Ultimately, the trajectory simulation gives a means for predicting 15 QoIs describing the flight state given a sample of 76 uncertain input parameters (i.e., $Z\in\mathbb{R}^{76}$ in Equation (4)) including those describing the deployment orientation and atmospheric conditions. Though the MC method was shown to produce reasonable results, the analysis was revisited with a multifidelity uncertainty propagation approach to significantly improve computational efficiency [3]. Four models were used in that work: (1) the high-fidelity POST2 model used in the original MC-based work, (2) a POST2 model with reduced physics, i.e., simplified atmospheric model, (3) a POST2 model with coarse timestep, i.e., reduced integration accuracy, and (4) a machine learning surrogate model where a support vector machine was trained on samples of the high-fidelity model. The models had varying levels of computational speedup compared to the high-fidelity model, as seen in Table 2. Here, it is assumed that the training of the machine learning surrogate model (including generation of training data) is an offline cost and is thus not factored into the speedup value. Notably, the reduced physics and coarse timestep also had hyperparameters ($\Delta t$) which were hand-selected with best engineering judgment. This work focuses on exploring the benefits of automated model tuning of the time step hyperparameters relative to the performance of using hand tuned values.
Table 2: Models considered for multifidelity uncertainty propagation.
@@
The proposed approach was implemented using Sandia National Laboratories’ uncertainty quantification code, Dakota. Figure 3 shows a diagram illustrating the implementation approach. Dakota controls the solution of the bi-level model tuning optimization problem (Equation (13)), using global surrogate-based optimization (i.e., the EGO algorithm) on the outer loop, passing values of $\underline{\beta}$ to ACV sample allocation optimization at each iteration. To solve the ACV problem, calls are dispatched to high- and low-fidelity models for pilot sampling to estimate the required model statistics for calculating estimator variance.
@@ 4 Results
This section presents results from two numerical examples that apply model tuning for multifidelity uncertainty propagation in the context of the ADEPT application. The first example focuses on tuning the time step of the Coarse Time Step model (Table 2) only. In this simpler 1D setting, the end-to-end, online uncertainty propagation process depicted in Figure 1 is carried out to evaluate the performance of the approach in a realistic setting where pilot sampling costs and errors are accounted for in the construction of multifidelity estimators.
@@
    Hand Selected - multifidelity estimators generated with fixed models borrowed from previous work [3] that were selected with engineering judgement (Table 2). For example 1, online pilot solution mode will be used to estimate model statistics on the fly during the sample allocation optimization.
@@
    Best Case - multifidelity estimators generated with a priori knowledge of both optimal hyperparameter values (no tuning cost) and model statistics (offline pilot solution mode with $N_{pilot}=10k$).
@@
Figure 5: Variance reduction as a function of low-fidelity model time steps for the QoIs considered in example 1 (a) and example 2 (b, c, d).
Figure 4 illustrates the effect of the tunable time step parameters on the low-fidelity models’ computational cost and correlations versus the other models for the time of flight QoI. While the cost of both the Reduced Physics and Coarse Time Step models decrease consistently with increasing time step as expected, the model correlations exhibit interesting and non-intuitive behavior. The correlations of both tunable models with respect to the High Fidelity and Machine Learning models exhibit relatively constant correlation over a large range of increasing time step, but show a drastic decrease/increase with respect to the High Fidelity/Machine Learning around $\Delta t=10^{-1}$ sec. The complex nature of this cost versus correlation trade-off makes it difficult to make a priori assumptions about valid hand tuned time step values. Note that ACV methods also take advantage of the correlations amongst the low-fidelity models themselves, further complicating manual hyperparameter tuning.
@@
that is possible with the ACV-MF estimator as a function of the tunable time step parameters for a computational budget of 1000 equivalent high-fidelity model evaluations. Here, each time step parameter was discretized as a grid of 25 equally (log) spaced points and the estimator variance, $\mathbb{V}ar\!\left[\tilde{Q}\right$], was evaluated at each point using offline pilot solution mode with $N_{pilot}=10k$. Example 1 is a 1D model tuning problem focused on estimating the time of flight QoI (Figure 5 (a)) while example 2 is a 2D model tuning problem considering the time of flight, landing latitude, and angle of attack at 10km QoIs (Figure 5 (b)-(d)). These plots help to illustrate the nature of the model tuning optimization problem that must be solved by providing a visualization of the objective function to be minimized with EGO. The plots also show the time step values for the Hand Selected and Best Case solutions, highlighting the large difference between those settings as well as the resulting variance reduction. In particular, the optimal time step setting for the Reduced Physics model is significantly larger than the hand selected value for all three QoIs in Figures 5 b-d. This indicates that there may be a tendency by practitioners to overestimate how accurate a low-fidelity model must be to be useful within a multifidelity framework, where models with large absolute errors can still be valuable provided they are adequately correlated.
@@ 4.1 Example 1 - Single Tunable Hyperparameter
Example 1 carries out the full online multifidelity estimator procedure with model tuning and iterated ACV shown in Figure 1, using the bilevel optimization strategy in Equation (13) to select an optimal timestep for the Coarse Time Step model. Estimators for the time of flight QoI are formed using different numbers of pilot samples, $N_{pilot}=\{10,50,100\}$, and EGO iterations, $N_{iter}=\{5,10,20\}$ [^8] EGO uses an initial sample size $N_{init}=3$ to generate a surrogate model, so these values represent 2, 7, and 17 EGO iterations beyond the initial sample, respectively, to solve the model tuning optimization problem. One hundred random trials are performed for each model tuning estimator as well as the Monte Carlo, Hand Selected, and Best Case solution types described above for three different computational budgets, $C_{budget}=\{500,1000,2000\}$, defined in terms of equivalent high-fidelity model cost. Note that the Best Case solution uses an optimal time step value of $\Delta t=0.065\;\mathrm{sec}$ and the Hand Selected case uses $\Delta t=0.1\;\mathrm{sec}$ (see Figure 5 a). Each solution approach is compared in terms of the MSE defined in Equation (16) with a reference solution computed using Monte Carlo with $N=100k$ samples. For all ACV solutions, the ACV-MF estimator (Table 1) is used.
Figure 6: Optimal time steps found across 100 random trials for example 1 for different numbers of a) pilot samples and b) EGO iterations.
@@
Results for the solution of the model tuning optimization problem (Equation (13)) are shown in Figures 6 and 7. Figure 6 shows the optimal time step values returned by the EGO algorithm across the 100 random trials for different numbers of $N_{pilot}$ and $N_{iter}$, represented as histograms. Compared with the variance reduction curve that is being maximized [^9] Note that maximizing the variance reduction is equivalent to minimizing the estimator variance which was the approach implemented in this work (Equation (13)), it can be seen in (a) that the EGO-optimized time steps noticeably improve with increasing $N_{pilot}$. In particular, for $N_{pilot}=10$, it can be seen that a relatively large percentage of optimized time steps are larger values that are sub-optimal in the sense that the variance reduction is low. For increasing $N_{iter}$ in (b), the trend is less clear, indicating that $N_{iter}=5$ seems to be sufficient and there are diminishing returns beyond this value. More detailed optimization results can be found in Appendix B and Figure 17. Note that while the variance reduction objective in Figure 6 appears simple to optimize, this curve was generated using $N_{pilot}=10k$ and is particularly noisy for the smaller values of $N_{pilot}$ used for model tuning, making the model tuning optimization problem challenging even in one dimension. This challenge is expanded upon in Appendix C and Figure 18. As noted in Section 2.5, the online pilot sampling approach could be used rather than pilot projection during model tuning optimization to yield more precise variance estimates, at the expense of adaptively evaluating additional pilot samples.
@@
Figure 7: The computational overhead ($\%$) from solving the model tuning optimization problem in example 1 for different numbers of a) pilot samples and b) EGO iterations.
@@
Figure 7 presents the computational cost of the model tuning optimization problem for different numbers of $N_{pilot}$ and $N_{iter}$. Here, the model tuning overhead is represented as a percentage of the total budget, $100\times\frac{C_{\mathcal{T}}}{C_{budget}}$, for different total budgets $C_{budget}=\{500,1000,2000\}$. Since there is variability in the progression of the EGO algorithm, the overheads are shown as box plots illustrating the variation across the 100 random trials. Intuitively, this shows that the overhead incurred by model tuning prior to multifidelity uncertainty propagation is less significant with increasing total computational budget. So for $C_{budget}=500$, using $N_{pilot}=100$ or $N_{iter}=20$ depletes a large percentage of the budget (between $10-30\%$), while these percentages drop below $5\%$ for $C_{budget}=2000$. Table 3 summarizes the performance obtained through model tuning optimization along with the incurred overhead cost. Here, the $\%$ of random model tuning trials that returned hyperparameters that result in lower estimator variance than the Hand Selected solution are reported along with the average tuning cost (in terms of number of high-fidelity simulations). It can be seen that performance increases with increasing $N_{pilot}$, albeit with diminishing returns, while performance does not increase with increasing $N_{iter}$, despite dedicating increased cost for model tuning. This result indicates that it is more worthwhile to consider larger values for $N_{pilot}$ rather than $N_{iter}$ for this one-dimensional tuning problem.
@@
Figure 8: MSE versus budget for model tuning ($N_{pilot}=50$, $N_{iter}=5$) versus hand-tuned models, best-case ACV, and Monte Carlo solutions.
@@
Figure 9: MSE versus number of pilot samples compared with the Monte Carlo, hand-selected, and best case estimators for different computational budgets.
Figure 10: MSE versus number of EGO iterations compared with the Monte Carlo, hand-selected, and best case estimators for different computational budgets.
@@ 4.2 Example 2 - Tuning of Multiple Hyperparameters
Example 2 studies the tuning of time steps for both the Reduced Physics, $\Delta t_{RP}$, and Coarse Time Step, $\Delta t_{CST}$, models through the solution of Equation (13). Similar to Example 1, model tuning performance is evaluated for different numbers of pilot samples, $N_{pilot}=\{10,50,100\}$, and EGO iterations, $N_{iter}=\{10,20,30\}$ [^10] EGO uses an initial sample size $N_{init}=6$ to generate a surrogate model in two dimensions, so these values represent 4, 14, and 24 EGO iterations beyond the initial sample, respectively. Performance was assessed in terms of estimator variance, rather than the more computationally-intensive MSE considered in Example 1, allowing for the consideration of additional QoIs (the landing latitude and angle of attack at 10km) and larger computational budgets. The variance reduction contours for each of the three QoIs can be seen in Figure 5 (b)-(d), showing the varied nature of the model tuning optimization problems (perhaps simpler for landing latitude relative to time of flight and angle of attack). A maximum computational budget of $C_{budget}=10000$ is considered to better understand the relative advantages of model tuning for more realistic analyses [^11] The original Monte Carlo analysis for ADEPT used $C_{budget}=8000$ high-fidelity simulations [31]. Note that while the model tuning procedure is simply repeated for each QoI individually for the results presented next, the approach could be extended to find hyperparameters that perform optimally for the joint estimation of all QoIs simultaneously using multioutput ACVs [42].
Figure 11: Model tuning optimization results across random trials for different numbers of pilot samples for a) time of flight, b) landing latitude, and c) angle of attack QoIs.
Figure 12: The computational overhead ($\%$) from solving the model tuning optimization problem in example 2 for different numbers of a) pilot samples and b) EGO iterations.
@@
The performance of the model tuning solutions was assessed with respect to the other baseline approaches in terms of variance reduction with respect to the Monte Carlo solution (Equation 17). In order to reduce the computation time needed to estimate the variance reduction, an interpolator, $\tilde{\mathbb{V}}$, was constructed to approximate $\mathbb{V}ar\!\left[\tilde{Q}\right$]
@@
Here, $\tilde{\mathbb{V}}$ is constructed from estimator variance data estimated on a $25\times 25$ grid of equally (log) spaced time steps using offline pilot solution mode with $N_{pilot}=10k$ (i.e., the same data used to produce the contour plots in Figure 5). For the Hand Selected solution, $\tilde{\mathbb{V}}$ was evaluated with $\Delta t_{RP}=0.001$ and $\Delta t_{CST}=0.1$ while the Best Case solutions used different optimal time step settings for the different QoIs considered (see Figure 5). For the model tuning solution, the estimated variance from the interpolator was scaled to account for the computational overhead of solving Equation (13),
@@
Figure 13: Model tuning performance for different numbers of pilot samples and computational budgets for a) time of flight, b) landing latitude, and c) angle of attack QoIs.
Figure 14: Model tuning performance for different numbers of EGO iterations and computational budgets for a) time of flight, b) landing latitude, and c) angle of attack QoIs.
@@
Figure 15: Model tuning performance when combined with parametric ACV search for different numbers of pilot samples for a) time of flight, b) landing latitude, and c) angle of attack QoIs.